OpenStudy (anonymous):

guys just a question i need to make sure of: for an angle (-pie/4) what is the cos, sin, tan, cot, sec, and csc plz

OpenStudy (nubeer):

sin(pi/4) = 1/(2)^0.5 cos(pi/4) = 1/(2)^0.5

OpenStudy (anonymous):

what?? no cos(pi/4)= (2)^0.5/2 and the sin is equal

OpenStudy (nubeer):

ohh its the same thing just rationalize what i have wrote it will give the same answer as urs. sin and cos pi/4 have same value.